one exterior angle of a regular polygon measures 10 degrees. what is the measure of each interior angle? How many sides does the polygon have ?
Answer by edjones(7569)

a) 180-10=170 degrees
.
b)
n=number of sides.
interior angle= (n-2) * 180/n = 170
(180n-360)/n=170
180n-360=170n
10n=360
n=36 sides
.
Ed